# Probing Pre-trained Language Models on Code Changes: Insights from ReDef, a High-Confidence Just-in-Time Defect Prediction Dataset

### v1

@@ 6.1. Internal Validity
Choice of Input Encoding. We evaluated five representative encodings, including formats inspired by prior JIT-SDP studies. However, given that all encodings ultimately derive from function_before and function_after, other ways of structuring or combining these views may provide richer signals of semantic change. For example, tree-structured or graph-based encodings could expose different aspects of program structure, and hybrid schemes may better capture interactions between edits and surrounding context. Exploring such alternatives remains an open avenue for future work.
Statistical Power. All statistical tests were based on 10 independent runs with different seeds. While this design controls stochastic variance, a larger number of repetitions could yield stronger statistical power and narrower confidence intervals. However, prior AI studies on code intelligence tasks rarely report extensive repetitions: many omit random seeds entirely or rely on a single run, and even the more rigorous works seldom exceed 10 repetitions due to the substantial computational cost of PLMs (Ye et al., 2023; Grishina et al., 2023). Running 30 or more trials per configuration is generally infeasible in practice. Our choice of 10 runs follows this common practice while still providing a reasonable trade-off between robustness and resource constraints.
@@ 6.2. External Validity
Project Selection Bias. Our dataset was drawn from large, long-lived infrastructure projects (e.g., Linux, FFmpeg, Chromium). These may differ from smaller applications or libraries in review culture, test maturity, and revert policies. We partially mitigated this risk by including diverse domains and reporting a leave-one-project-out analysis, but broader coverage is needed. In particular, industrial or proprietary projects could exhibit markedly different defect patterns, and including such settings would provide stronger external validity.
@@ 7. Conclusion
This work introduced ReDef, a high-confidence JIT-SDP dataset of function-level modifications with both defective and clean cases: 3,164 defective and 10,268 clean function-level modifications, with manual audits confirming 92% label precision—substantially higher than the prior JIT-SDP corpora. Our curation pipeline combines revert anchoring, longitudinal history screening, and GPT-assisted triage with multiple votes and audits to reduce label noise. Unlike conventional SZZ-based resources which infer bug-inducing commits post hoc and often lack clean counterparts, ReDef offers the community a reliable benchmark for studying change-aware models, enabling reproducible comparison and facilitating exploration of edit-sensitive learning objectives.
Using ReDef, we fine-tuned three strong PLMs—CodeBERT, CodeT5+, and UniXcoder—under five input encodings. Compact, change-focused encodings and tagged diffs achieved the best results, whereas whole-function inputs often diluted the signal and suffered from truncation under token limits. Notably, UniXcoder performed worse at a 1,024-token budget than at 512, showing that more context does not necessarily help when change information becomes obscured.
Counterfactual perturbations of change representations yielded almost no performance degradation. Even when we swapped the order of deleted and added blocks or mislabeled unchanged lines as modified, F1 remained nearly identical. If PLMs had truly internalized the semantics of edits, such distortions should have impaired performance. This outcome exposes a critical blind spot in how current PLMs are evaluated on code modifications: what appears to be robustness is, in fact, an artifact of surface-level cue reliance. Our findings caution against interpreting strong benchmark scores as evidence of genuine change understanding.
Threats to validity remain. We examined only five encoding strategies, but alternative ways of presenting function_before and function_after may provide richer semantic signals. Our dataset focuses on large C/C++ projects, leaving open questions about smaller repositories and other languages. Finally, although we repeated all experiments with ten random seeds, stronger statistical power could be obtained with more runs.
Future work should probe why PLMs lag in modification understanding compared to single-snapshot tasks such as vulnerability detection. Promising directions include pre-training objectives that explicitly expose models to edits and diffs, as well as richer change representations or architectures that make modification semantics more accessible to PLMs. Beyond PLMs, evaluating decoder-based code LLMs (e.g., CodeLlama, StarCoder) and API models (e.g., GPT-4o) under the same protocol will test whether larger generative models can close this gap.

### v2

@@ 6.1. Internal Validity
Subjectivity in Manual Verification. Determining whether a code modification is actually bug-inducing is an inherently difficult and subjective task. To mitigate this subjectivity, our manual audit process relied primarily on explicit commit-message evidence (e.g., statements such as “reverting due to a crash”) rather than a subjective interpretation of code logic. This is because such natural-language mentions provide the most direct and reliable justification from the developers who performed the revert. In cases where the commit messages were vague or lacked sufficient reasoning for the revert, we categorized them as incorrectly collected samples. Consequently, this conservative auditing approach revealed an error rate of only 8% (Table 3), thereby demonstrating the high precision of our curated dataset. Despite these rigorous criteria, human inspection is not infallible, and the potential for misjudgment during the verification of the 100 samples remains a threat to internal validity.
Input Encodings and Perturbation Strategies. We evaluated five representative encodings, including formats inspired by prior JIT-SDP studies. However, all encodings in our study are derived solely from function_before and function_after, which prevents us from leveraging richer semantic signals available in other sources such as commit messages. Moreover, even within these code-centric views, alternative ways of structuring the information—such as tree-structured or graph-based encodings—may provide richer signals of semantic change that are not fully captured by the linear, token-based representations used in this work. Additionally, we designed four perturbations as diagnostic probes specifically intended to assess semantic sensitivity. While our results demonstrated that models remain largely indifferent even under significant semantic disruptions, there may exist alternative perturbation strategies that could further expose the extent of this insensitivity or uncover different facets of modification reasoning. Exploring such alternatives and their alignment with diverse real-world scenarios remains an open avenue for future work.
Statistical Power. All statistical tests were based on 10 independent runs with different seeds. While this design controls stochastic variance, a larger number of repetitions could yield stronger statistical power and narrower confidence intervals. However, prior AI studies on code intelligence tasks rarely report extensive repetitions: many omit random seeds entirely or rely on a single run, and even the more rigorous works seldom exceed 10 repetitions due to the substantial computational cost of CLMs (Ye et al., 2023; Grishina et al., 2023). Running 30 or more trials per configuration is generally infeasible in practice. Our choice of 10 runs follows this common practice while still providing a reasonable trade-off between robustness and resource constraints.
@@ 6.2. External Validity
Dataset Coverage and Generalizability. We intentionally restricted the scope of ReDef to single-function modifications anchored by revert commits. Since reverts and single-function edits represent only a subset of all defective modifications, our dataset does not encompass the full spectrum of defect patterns, such as multi-function changes or bugs that are not easily identified for reversion. This focus may limit the generalizability of our findings to more complex, “tangled” commits. However, this conservative approach significantly mitigates the “blame-innocent” problem inherent in traditional SZZ-based labeling, ensuring high label precision. By prioritizing reliability over sheer volume, ReDef serves as a high-confidence benchmark that complements existing large-scale but noisy datasets, providing a necessary trade-off between coverage and predictive reliability.
Project Selection Bias. Our dataset was drawn from 22 large-scale open-source projects across diverse domains. However, the data volume is notably concentrated in a few massive repositories, particularly Linux and Chromium. This skew risks biasing our findings toward system-level software patterns, potentially limiting their generalizability to smaller applications or projects with different review and revert policies. Furthermore, industrial or proprietary projects could exhibit markedly different defect patterns, and including such settings in future work would provide stronger external validity.
@@ 7. Conclusion
This work introduced ReDef, a high-confidence JIT-SDP dataset consisting of 3,164 defective and 10,268 clean function-level modifications. Manual audits confirmed a label precision of 92%, which is substantially higher than that of prior JIT-SDP corpora. Our curation pipeline—combining revert anchoring, longitudinal history screening, and GPT-assisted triage—effectively minimizes the label noise prevalent in existing datasets. By overcoming the limitations of conventional SZZ-based resources, ReDef provides a robust foundation for the research community to evaluate change-aware models. This benchmark not only enables reproducible comparisons but also facilitates the development of next-generation, edit-sensitive learning objectives.
Using ReDef, we evaluated four representative Code Language Models (CLMs)—CodeBERT, CodeT5+, UniXcoder, and Qwen2.5—across five input encoding strategies. Our results indicate that compact, change-focused encodings consistently achieve the best performance by emphasizing localized modifications and minimizing information loss due to truncation. Notably, we observed that larger context windows do not necessarily translate to better performance; both UniXcoder and Qwen2.5 performed worse under a 1,024-token budget than at 512 tokens, suggesting that excessive context can obscure critical change signals.
Counterfactual perturbations of change representations further revealed that CLMs’ performance remains effectively stable even under significant structural distortions. While encoder-based models were almost entirely unaffected by these perturbations, the decoder-based Qwen2.5 exhibited only marginal sensitivity in specific conditions. Even when we reversed the logical order of modifications or injected spurious change markers, F1 remained nearly identical. If these models had truly internalized the semantics of code edits, such fundamental distortions should have led to a precipitous decline in predictive performance.
This outcome exposes a critical blind spot in current evaluations of code-change models: what appears to be robustness is, in fact, a symptom of semantic blindness, where models fall back on surface-level cues rather than genuine understanding. Our findings caution against interpreting high benchmark scores as evidence of an advanced grasp of code evolution. Ultimately, this study underscores the need for new architectures and training objectives specifically designed to capture the relational and causal dynamics of code transformations.
Several threats to validity remain. While ReDef ensures high label precision, its focus on single-function reverts in large-scale C/C++ projects may limit generalizability to multi-function commits or other software domains. Furthermore, despite our manual audits, the inherent subjectivity of human inspection persists as a potential bias, and alternative encoding or perturbation strategies may exist that could uncover deeper semantic insights. Future research should explore the underlying causes of this semantic lag, potentially through edit-aware pre-training objectives or architectures specifically tailored for change reasoning. It is equally vital to investigate whether this reliance on superficial patterns persists in other semantics-sensitive tasks, such as vulnerability detection and automated code review. Finally, evaluating larger API-based models (e.g., GPT-4o) under the same protocol will reveal whether increased scale can effectively bridge the critical “semantic gap” exposed in this study.